# Rotary Masked Autoencoders are Versatile Learners

### v1

@@ Abstract
Applying Transformers to irregular time-series typically requires specializations to their baseline architecture, which can result in additional computational overhead and increased method complexity. We present the Rotary Masked Autoencoder (RoMAE), which utilizes the popular Rotary Positional Embedding (RoPE) method for continuous positions. RoMAE is an extension to the Masked Autoencoder (MAE) that enables representation learning with multidimensional continuous positional information while avoiding any time-series-specific architectural specializations. We showcase RoMAE’s performance on a variety of modalities including irregular and multivariate time-series, images, and audio, demonstrating that RoMAE surpasses specialized time-series architectures on difficult datasets such as the DESC ELAsTiCC Challenge while maintaining MAE’s usual performance across other modalities. In addition, we investigate RoMAE’s ability to reconstruct the embedded continuous positions, demonstrating that including learned embeddings in the input sequence breaks RoPE’s relative position property.
@@ 1 Introduction
Various specializations to the Transformer have been proposed to overcome this limitation. These can be divided into two main types: modifications of the internal architecture of the Transformer (e.g. modifying the feedforward layer in the Transformer Encoder Block [19, 47] and novel positional embeddings (e.g. using a neural ODE  [12]). Alternatively, modern State Space Models like Mamba [22] or S5 [51] are natively able to model various modalities such as text and images in addition to irregular time-series. Extending the capability of Transformers to irregular time-series while staying within existing frameworks developed for fields such as Natural Language Processing (NLP) and computer vision would allow one to easily benefit from ongoing developments within the Transformer “ecosystem”.
@@
    Continuous Positional Embedding with RoPE We investigate how RoPE can be used to embed continuous positions, expanding on its original concept introduced in the RoFormer architecture [54], which did not address non-uniform or real-valued timestamps. We show that learned input embeddings can invalidate RoPE’s relative distance guarantees, and we provide new empirical evidence that the information encoded by RoPE does not decay as positions grow further apart.
@@
This work is structured as follows: Section 2 covers related works on MAE, RoPE, and irregular time-series. Section 3 provides necessary background details. Section 4 details the workings of RoMAE and establishes the theoretical framework we use to tackle a variety of modalities. Section 5 presents the results of our experiments. Section 6 discusses our results. Section 8 concludes the paper.
Figure 1: Overview of the RoMAE pipeline. Left: Visualisation of data embedding via multi-dimensional (ND) patchification for demonstrative data realisations in 1, 2 and 3D. Centre: Full depiction of RoMAE  architecture. The optional [CLS] token is omitted from the input sequence for simplicity. Right: The RoMAE encoder/decoder with ROPE operations denoted by rotational arrows.
@@ 3 Background
Attention: We follow the formulation for Attention proposed in the original Transformer [59]. Let $\mathbf{z}$ be a sequence of $k$ embeddings $z_{i}\in\mathbb{R}^{d_{\text{model}}}$ for $i\in[1,2,...,k]$ and $d_{\text{model}}$ be the dimensionality of the model. Three linear layers are applied to transform $\mathbf{z}$ into sequences containing queries ($q_{i}$), keys ($k_{i}$), and values ($v_{i}$). E.g., $q_{i}=W_{q}z_{i}$$k_{i}=W_{k}z_{i}$$v_{i}=W_{v}z_{i}$. The matrices $Q,K,V$ containing $q$,$k$,$v$ are then passed through Scaled Dot-Product Attention (SDPA) as defined in Equation (1):
$\text{Attention}(Q,K,V)=\text{softmax}\left(\frac{QK^{T}}{\sqrt{d_{k}}}\right)V$ (1)
@@
Regular and Irregular Dimensions: RoMAE is designed to work with both regular and irregular dimensional data. Specifically, we consider inputs of the form: $\mathbf{x}\in\mathbb{R}^{d_{1}\times d_{2}\times\cdots,\times d_{D}}$ where $D$ is the number of dimensions in $\mathbf{x}$ and $d_{i}$ is the size of dimension $i$ for $i\in[1,\cdots,D]$.
@@
A regular dimension is one where all points are equally spaced, while an irregular dimension is one where the space between points varies.
Some dimensions in the input $\mathbf{x}$ may be irregular while others could be regular. For example, a set of irregularly sampled images has height and width as two regular dimensions and time as one irregular dimension.
@@ 3.1 Rotary Positional Embeddings
$p$-RoPE: In this work we make use of $p$-RoPE [4], a truncated version of RoPE where only the smallest $p$ percent of $\theta_{i}$’s are kept. This cuts out a fraction $(1-p)$ of the frequencies in RoPE and thus leaves a portion of the input embedding space unchanged. The unchanged region in the embedding space provides the model with a data channel it can use to pass information into SDPA without any modifications by RoPE, making the model more robust to varying sequence length. We use a value of 0.75 for $p$, which has been shown by Barbero et. al. [4] to work well.
@@ 4 Method
An overview of RoMAE can be seen in Figure 1.
N-Dimensional (ND) Patchification: Given inputs $\mathbf{x}$ as described in Section 3, we define a patch size $(p_{1},\cdots,p_{D})$ and divide each dimension into $N_{i}=d_{i}/p_{i}$ non-overlapping segments where $d_{i}$ is the size of dimension $i$. These are flattened, creating a sequence of patches with length $k=\prod_{i=1}^{D}N_{i}$ and number of elements per patch $n_{p}=\prod_{i=1}^{D}p_{i}$. Finally, a linear layer $W^{n_{p}\times d_{\text{model}}}$ is applied to each patch to project it into the embedding dimension $d_{\text{model}}$. This process is illustrated in Figure 1 and is the same as the one employed by ViT [15] for images and ViViT [1] for video. It can also be understood as using non-overlapping $N$-dimensional convolutions with the number of channels equal to $d_{\text{model}}$. After ND-patchification, we have a sequence of embeddings $\mathbf{z}$, which can be passed into the RoMAE Encoder.
@@ 4.1 Overall Structure
RoMAE’s structure follows MAE’s, using an asymmetric encoder/decoder, with the encoder being much larger than the decoder. Although both the encoder and decoder in RoMAE are a Transformer Encoder similar to BERT [14], we bring over recent developments in NLP from models such as Llama [42]. Specifically, we utilize the popular Sigmoid Linear Unit [24, 16] for the non-linearity. We also use RMSNorm [66] instead of Layer Normalization [2], as it has been shown to be more computationally efficient while maintaining the same level of model convergence. For more in-depth architectural details, including definitions of various model sizes, we refer to Appendix A.1.
@@ 4.2 Positional Information in RoMAE
Dealing With Large Amounts of Irregular Dimensions: Although we are able to encode multi-dimensional positional information using Axial RoPE, this does not scale well to a large number of dimensions due to having to divide $d_{\text{model}}$ by $D$. To overcome this, we optionally reserve a dimension in Axial RoPE that is used to store the dimensional index $i$ for $i\in[1,2,\cdots,D]$. E.g., if an embedding belongs to dimension 4, it will receive a positional encoding of 4.
We use this method primarily when working with multi-variate time-series where each variate is irregular. When using this approach we also include the learned [CLS] token, which allows the model to know what dimension each sample belongs to despite RoPE being relative as a consequence of Proposition 4.2. Note that this technique is not required for irregular time-series where every variate is present at each time-step. In this case, all variates can be combined into 1 embedding by using an appropriate patch size.
@@ 4.3 Effects of Relative Position in RoMAE
Intuition: The [CLS] token provides the model with an "anchor". This allows the model to compare each input embedding to the [CLS] token, and reconstruct its absolute position. We provide a proof in Appendix B, as well as experimental evidence in Section 5.1.
@@
Discussion: A claimed property of RoPE is that it causes the dot product between queries and keys to decay as their positions grow further apart [54]. In Appendix A.3, we provide empirical evidence showing that RoMAE is able to reconstruct position almost perfectly for a variety of distances. We also refer to the proof by Barbero et. al. [4], showing that it is possible to construct a key for any non-zero query and any distance such that the softmax value in Attention is maximized at that distance.
@@ 5 Experiments
Compute Details: The experiment on the DESC ELAsTICC Challenge [^1] https://portal.nersc.gov/cfs/lsst/DESC_TD_PUBLIC/ELASTICC/ (Section 5.3), was run on a Slurm cluster using 4 nodes for  4 hours with each having 4 Nvidia H100 (with 64GB HBM2 memory) GPUs.
The experiment on the Tiny Imagenet data set was run on one node of a Slurm cluster, utilizing two NVIDIA Tesla V100 GPUs for 5 hours.
Together, the experiments on the UEA Time Series Archive [3] (Section 5.3), Pendulum dataset [5] (Section 5.3), and absolute position experiments (Section 5.1) were run on a 1080ti GPU for a total of  1.5 hours.
@@ 5.1 Reconstructing Absolute Position
RoMAE-tiny 0.062 ($\pm$0.007) 200.33 ($\pm$0.001)
RoMAE-small 0.0057 ($\pm$0.002) 200.33 ($\pm$0.001)
RoMAE-base 0.0031 ($\pm$0.002) 200.33 ($\pm$0.002)
@@
We observe a clear difference between the model that uses the [CLS] token and the one that does not. When supplied with the learnable token, RoMAE is able to reconstruct the original absolute position almost perfectly. Larger models seem to achieve a better MSE, although all sizes perform well. For all experimental details and an additional experiment on absolute position reconstruction we refer to Appendix C.5 and Appendix A.3 respectively.
@@ 5.2 Tiny ImageNet
Table 2: Results Tiny ImageNet across various versions of RoMAE-small
Model precision recall F-score
RoMAE ([CLS]) 0.3427 0.3582 0.3418
RoMAE (no [CLS]) 0.3450 0.3662 0.3471
RoMAE (absolute) 0.3424 0.3638 0.3424
To investigate the effect of positional embedding and the learned [CLS] token on RoMAE, we train three versions of RoMAE on Tiny ImageNet [31]; RoPE with the [CLS] token, RoPE without the [CLS] token, and absolute sinusoidal positional embeddings [58] with the [CLS] token. When fine-tuning RoMAE without the [CLS] token, we place the classification head on top of the mean of the output embeddings.
The final configuration with absolute positional embeddings is very similar to MAE, making for a good comparison. We use a patch size of (16, 16) and mask 75% of the input, similar to MAE. After pre-training each model for 200 epochs, we fine-tune for another 15. For full experimental details we refer to Appendix C.1. The results are shown in Table 2. Although all 3 models perform similarly, the models using RoPE required a different learning rate to the model with absolute position to converge. We also highlight that although RoMAE with RoPE and no [CLS] token is invariant to translation (Proposition 4.1), this does not seem to affect the scores on images. Overall however RoMAE seems to perform as well as MAE on images. As an additional experiment comparing RoMAE to standard MAE, we train RoMAE to classify audio files after a self-supervised pretraining on unlabeled audio data. This experiment is illustrated in Appendix A.4.
@@ 5.3 Irregular Time-series Classification
Transformer [58, 7] ATAT [7] RoMAE-tiny
F-score 0.5256 0.6270 0.6770
The DESC ELAsTiCC Challenge is a multi-variate irregular time-series dataset consisting of 1,845,146 simulated light curves and 36 classes of astronomical objects. Each light curve consists of 6 irregularly sampled channels. To embed this data in RoMAE we follow the procedure described in Section 4.2. Although ELAsTiCC has additional metadata for each light curve, we compare performance only across raw light curves. We train RoMAE-small by conducting full pre-training for 200 epochs with a masking ratio of 75%, then fine-tuning for 25 epochs. For all details regarding ELAsTiCC experimental setup and a visualization of the data we refer to Appendix C.6.
Table 3 shows our results using RoMAE, and compares with ATAT [7], a Transformer architecture specialized for ELAsTiCC. Despite the latter’s specialization, RoMAE improves over ATAT by about .05 F-score. One key difference between ATAT and RoMAE is that ATAT does not conduct any pre-training, which may be a key reason for RoMAE’s better performance.
UEA Multivariate Time-series Archive
Model
Dataset TST [65] mTAN [45] S5 [51] ContiFormer [11] RoMAE
BM 0.9667 0.9917 0.9833 0.9750 0.9917
CT 0.9742 0.9529 0.9610 0.9833 0.9882
EP 0.9589 0.9203 0.9074 0.9324 0.9517
HB 0.7398 0.7789 0.7333 0.7561 0.7447
LSST 0.5520 0.5307 0.6389 0.6004 0.6225
We evaluate RoMAE on a variety of datasets from the UEA Multivariate Time-series Archive [3]. To make the datasets irregular we follow the procedure by Kidger et. al. [30], dropping 30% of the observations. Because all variates are present at each time-step, the only irregular dimension is time. Therefore, to embed this data in RoMAE, we combine all variates per time-step into one embedding. This is a much simpler setup than the one used in Section 5.3. For each dataset we conduct pre-training for 800 epochs. When fine-tuning, we found it necessary to change hyper-parameters between different datasets. For more details on our experimental setup we refer to Appendix C.2.
Table 4 presents the mean accuracy from 3 full training runs (pretraining + finetuning), as well as published results from a variety of models. Overall, RoMAE is as good or better than the comparators, and is able to handle the various datasets without issues. All the datasets trained on are relatively small, with some being on the order of hundreds of samples. Therefore, this experiment also shows how data-efficient RoMAE can be.
Irregular Time-series Regression: Pendulum Dataset
Table 5: Regression MSE ×$\times 10^{-3}$ (mean $\pm$ std) on the Pendulum dataset.
Model Regression MSE (×$\times 10^{-3}$)
@@
The Pendulum dataset [51] is an irregular time-series dataset consisting of images of a pendulum. To embed the images in RoMAE, we use a patch size of $(1,24,24)$ for (time, height, width). This corresponds to 1 embedding per time-step/image. RoMAE is trained directly on regression without any pre-training, predicting the sine and cosine of the angle of the pendulum which follows a non-linear dynamical system. To stay close in size to the other models evaluated on this dataset, we instantiate RoMAE with only 2 layers and an MLP hidden size of 30. For additional information on the dataset and full experimental details we refer to Appendix C.3. After training on 20 different seeds, the mean MSE and standard deviation of RoMAE on the Pendulum dataset are reported in Table 5. RoMAE is able to achieve an extremely competitive score, outperforming specialized Transformer based models such as ContiFormer [12], as well as state space models such as S5 [51].

### v2

@@ Abstract
Applying Transformers to irregular time-series typically requires specializations to their baseline architecture, which can result in additional computational overhead and increased method complexity. We present the Rotary Masked Autoencoder (RoMAE), which utilizes the popular Rotary Positional Embedding (RoPE) method for continuous positions. RoMAE is an extension to the Masked Autoencoder (MAE) that enables interpolation and representation learning with multidimensional continuous positional information while avoiding any time-series-specific architectural specializations. We showcase RoMAE’s performance on a variety of modalities including irregular and multivariate time-series, images, and audio, demonstrating that RoMAE surpasses specialized time-series architectures on difficult datasets such as the DESC ELAsTiCC Challenge while maintaining MAE’s usual performance across other modalities. In addition, we investigate RoMAE’s ability to reconstruct the embedded continuous positions, demonstrating that including learned embeddings in the input sequence breaks RoPE’s relative position property.
@@ 1 Introduction
Various specializations to the Transformer have been proposed to overcome this limitation. These can be divided into two main types: modifications of the internal architecture of the Transformer (e.g. modifying the feedforward layer in the Transformer Encoder Block [19, 46]) and novel positional embeddings (e.g. using a neural ODE  [12]). Alternatively, modern State Space Models like Mamba [22] or S5 [51] are natively able to model various modalities such as text and images in addition to irregular time-series. Extending the capability of Transformers to irregular time-series while staying within existing frameworks developed for fields such as Natural Language Processing (NLP) and computer vision would allow one to easily benefit from ongoing developments within the Transformer “ecosystem”.
@@
    Continuous Positional Embedding with RoPE: We investigate how RoPE can be used to embed continuous positions, expanding on its original concept introduced in the RoFormer architecture [54], which did not address non-uniform or real‐valued timestamps. We show that learned input embeddings can invalidate RoPE’s relative distance guarantees, and we provide new empirical insights into RoPE’s ability to embed varying positional scales.
@@
This work is structured as follows: Section 2 covers related works on MAE, RoPE, and irregular time-series. Section 3 provides the necessary background material. Section 4 details the workings of RoMAE and establishes the theoretical framework we use to tackle a variety of modalities. Section 5 presents the results of our experiments. Section 6 discusses our results. Section 7 concludes the paper.
Figure 1: Overview of the RoMAE pipeline. Left: Visualisation of data embedding via multi-dimensional (ND) patchification for illustrative data realisations in 1, 2 and 3D. Centre: Full depiction of RoMAE  architecture. The optional [CLS] token is omitted from the input sequence for simplicity. Right: The RoMAE encoder/decoder with ROPE operations denoted by rotational arrows.
@@ 3 Background
Attention: We follow the formulation for Attention proposed in the original Transformer [58]. Let $\mathbf{z}$ be a sequence of $k$ embeddings $z_{i}\in\mathbb{R}^{d_{\text{model}}}$ for $i\in[1,2,...,k]$ and $d_{\text{model}}$ be the dimensionality of the model. Three linear layers are applied to transform $\mathbf{z}$ into sequences containing queries ($q_{i}$), keys ($k_{i}$), and values ($v_{i}$). E.g., $q_{i}=W_{q}z_{i}$, $k_{i}=W_{k}z_{i}$, $v_{i}=W_{v}z_{i}$. The matrices $Q,K,V$ containing $q$,$k$,$v$ are then passed through Scaled Dot-Product Attention (SDPA) as defined in Equation (1):
$\text{Attention}(Q,K,V)=\text{softmax}\left(\frac{QK^{T}}{\sqrt{d_{\text{model}}}}\right)V$ (1)
@@
Regular and Irregular Dimensions: RoMAE is designed to work with both regular and irregular dimensional data. Specifically, we consider inputs of the form: $\mathbf{x}\in\mathbb{R}^{d_{1}\times d_{2}\times\cdots\times d_{D}}$ where $D$ is the number of dimensions in $\mathbf{x}$ and $d_{i}$ is the size of dimension $i$ for $i\in[1,\cdots,D]$.
@@
A regular dimension is one where all points are equally spaced, while an irregular dimension is one where the distance between points varies.
Some dimensions in the input $\mathbf{x}$ may be irregular while others could be regular. For example, a set of images sampled at irregular times has height and width as two regular dimensions and time as one irregular dimension.
@@ 3.1 Rotary Positional Embeddings
$p$-RoPE: In this work we make use of $p$-RoPE [4], a truncated version of RoPE where only the $p$ percent of smallest $\theta_{i}$ values are kept. This cuts out a fraction (1p)(1-p) of the frequencies in RoPE and thus leaves a portion of the input embedding space unchanged. The unchanged region in the embedding space provides the model with a data channel it can use to pass information into SDPA without any modifications by RoPE, making the model more robust to varying sequence length. We use a value $p=0.75$, which has been shown by Barbero et. al. [4] to work well.
@@ 4 Method
An overview of RoMAE is shown in Figure 1.
N-Dimensional (ND) Patchification: Given inputs $\mathbf{x}$ as described in Section 3, we define a patch size $(p_{1},\cdots,p_{D})$ and divide each dimension into $N_{i}=d_{i}/p_{i}$ non-overlapping segments, where $d_{i}$ is the size of dimension $i$. These are flattened, creating a sequence of patches with length $k=\prod_{i=1}^{D}N_{i}$ and number of elements per patch $n_{p}=\prod_{i=1}^{D}p_{i}$. Finally, a linear layer $W^{n_{p}\times d_{\text{model}}}$ is applied to each patch to project it into the embedding dimension $d_{\text{model}}$. This process is illustrated in Figure 1 and is the same as employed by ViT [15] for images and ViViT [1] for video. It can also be understood as using non-overlapping $N$-dimensional convolutions with the number of channels equal to $d_{\text{model}}$. After ND-patchification, we have a sequence of embeddings $\mathbf{z}$, which can be passed into the RoMAE Encoder.
@@ 4.1 Overall Structure
RoMAE’s structure follows MAE’s, using an asymmetric encoder/decoder, with the encoder being much larger than the decoder. Although both the encoder and decoder in RoMAE are a Transformer Encoder similar to BERT [14], we bring over recent developments in NLP from models such as Llama [43]. Specifically, we utilize the popular Sigmoid Linear Unit [24, 16] for the non-linearity. We also use RMSNorm [66] instead of Layer Normalization [2], as it has been shown to be more computationally efficient while maintaining the same level of model convergence. For architectural details, including definitions of various model sizes, we refer to Appendix A.2.
@@ 4.2 Positional Information in RoMAE
Dealing With Many Irregular Dimensions: Although we are able to encode multi-dimensional positional information using Axial RoPE, this does not scale well to a large number of dimensions due to having to divide $d_{\text{model}}$ by $D$. To overcome this, we optionally reserve a dimension in Axial RoPE that is used to store the dimensional index $i$ for $i\in[1,2,\cdots,D]$. E.g., if an embedding belongs to dimension 4, it will receive a positional encoding of 4. When using this approach we include the learned [CLS] token, which allows the model to recover the dimensional index despite RoPE being a RPE method (as a consequence of Proposition 4.2). We utilize this method with the ELAsTiCC dataset in Section 5.4 which allows us to reduce the number of positional dimensions from 6 to 2.
@@ 4.3 Effects of Relative Position in RoMAE
Intuition: The [CLS] token provides the model with an "anchor". This allows the model to compare each input embedding to the [CLS] token, and reconstruct its absolute position. We provide a proof in Appendix C, as well as experimental evidence in Section 5.1.
@@
Discussion: A claimed property of RoPE is that it causes the dot product between queries and keys to decay as their positions grow further apart [54]. In Appendix B.1, we provide empirical evidence showing that RoMAE is able to reconstruct position almost perfectly over two different scales of distance. We also refer to the proof by Barbero et. al. [4], showing that it is possible to construct a key for any non-zero query and any distance such that the softmax value in Attention is maximized at that distance.
@@ 5 Experiments
Compute Details:
The experiment on the Tiny ImageNet data set (Section 5.2) was run on one node of a Slurm cluster, utilizing two NVIDIA Tesla V100 GPUs for 5 hours.
The experiment on the DESC ELAsTICC Challenge [^1] https://portal.nersc.gov/cfs/lsst/DESC_TD_PUBLIC/ELASTICC/ (Section 5.4), was run on a Slurm cluster using 4 nodes for $\sim 4$ hours with each having 4 Nvidia A100 (with 64GB memory) GPUs.
Together, the experiments on the UEA Time Series Archive [3] (Section 5.4), Pendulum dataset [5] (Section 5.4), and absolute position experiments (Section 5.1) were run on a 1080ti GPU for a total of $\sim 1.5$ hours.
@@
Model and experimental code for RoMAE is made public through a convenient Python package. [^2] https://chromeilion.github.io/RoMAE-Website/
@@ 5.1 Reconstructing Absolute Position
RoMAE-tiny 0.062 (0.007) 200.33 (0.001)
RoMAE-small 0.0057 (0.002) 200.33 (0.001)
RoMAE-base 0.0031 (0.002) 200.33 (0.002)
Table 2: Results on Tiny ImageNet across various versions of RoMAE-small
Model F-score ($\pm$ std)
RoMAE (no [CLS]) 0.500 (0.011)
RoMAE ([CLS]) 0.475 (0.006)
RoMAE (absolute) 0.479 (0.010)
@@
We observe a clear difference between the model that uses the [CLS] token and the one that does not. When supplied with the learnable token, RoMAE is able to reconstruct the original absolute position almost perfectly. Larger models seem to achieve a better MSE, although all sizes perform well. For all experimental details and an additional experiment on absolute position reconstruction we refer to Appendix D.5 and Appendix B.1 respectively.
@@ 5.2 Tiny ImageNet
To investigate the effect of positional embedding and the learned [CLS] token on RoMAE, we train three versions of RoMAE on Tiny ImageNet [31]; RoPE with the [CLS] token, RoPE without the [CLS] token, and absolute sinusoidal positional embeddings [58] with the [CLS] token. When fine-tuning RoMAE without the [CLS] token, we place the classification head on top of the mean of the output embeddings, otherwise we place the head on top of the [CLS] token. The final configuration with absolute positional embeddings is very similar to MAE, making for a good comparison. We use a patch size of (16, 16) and mask 75% of the input, similar to MAE. After pre-training each model for 200 epochs, we fine-tune for another 15. We also follow the procedure outlined by MAE to compute the pre-training loss using normalized patch values instead of pixel values.
For full experimental details we refer to Appendix D.1. The results are shown in Table 2. Although all 3 models perform similarly, we highlight that RoMAE with RoPE and no [CLS] token performs slightly better than both RoPE with [CLS] and absolute positional embeddings. This could be because of the models translation invariance (Proposition 4.1). The difference could also come from placing the classification head on top of the mean of output embeddings instead of the [CLS] token. Overall, our results indicate that RoMAE performs at least as well as MAE on images.
5.3 Audio benchmark
Table 3: Results for the ESC-50 benchmark for audio datasets, for the RoMAE-small model.
Model Accuracy
SSAST (Librispeech) 80.0
RoMAE-small (Librispeech) 83.2
SSAST (AudioSet-20k) 82.2
RoMAE-small (AudioSet-20k) 84.7
We chose to test RoMAE’s ability to classify audio files, after a self-supervised pre-training on unlabeled audio datasets, inspired by the SSAST pretraining strategy [21]. SSAST is the first Vision transformer-based model that introduces a self-supervised pretraining strategy, supporting arbitrary patch size, for the audio representation learning.
In order to apply our model to the audio datasets, we first transform the audio waveforms to Mel spectrograms. First, the input audio waveform of length $t$ seconds is converted into a sequence of 128-dimensional log-Mel filterbank [^4] The Mel Filterbank transform computes weighted averages of bins to provide spectral power estimates on a logarithmic frequency scale, which is more affine to human hearing resolution. (fbank) features computed with a 25ms Hamming window every 10ms. This results in a 128 × 100 $t$ spectrogram. For the pretraining step, we followed the patchification strategy adopted in [21] and we split the spectrogram into a sequence of NN $(16\times 16)$ patches, where N=12(100t16)/10N=12(100t-16)/10 is the number of patches and the effective input sequence length for the model. We refer to Appendix D.2 for the list of hyperparameters adopted for the pretraining and finetuning of the model. We would like to highlight here that the pretraining was run for 150 epochs, without the [CLS] token, while we choose to adopt a mask ratio equal to 0.75, that is very similar to the masking fraction associated with the SSAST-patch 400 model.
For the finetuning audio classification benchmark, we used the ESC-50 dataset [39], consisting of 2000 5-second environmental audio recordings classified into 50 classes. The current best results on ESC-50 are accuracies of 86.5 and 94.7 obtained with supervised training (on AudioSet-2M) respectively by SOTA-S and SOTA-P models. The SSAST model, which is the only ViT model adopting self-supervised training for this task, achieved accuracies of 82.2 and 84.6 when trained, respectively, on AudioSet-20K and AudioSet-2M (as reported in Table 2 of [21]), showing that the size and richness of the pretraining dataset impacts, in a non-negligible way, the performance of that model on the finetuning tasks. Since we did not have sufficient computational resources to pretrain RoMAE on AudioSet-2M, we chose to compare our model performance with that of the SSAST model pretrained on AudioSet-20K and Librispeech, respectively. We report the benchmark results in Table 3, showing that RoMAE performs better than the SSAST model in these two cases.
@@ 5.4 Irregular Time-series Classification
Method F-score
Transformer [58, 7] 0.5256
ATAT [7] 0.6270
RoMAE-tiny-shallow 0.7106
RoMAE-tiny 0.8029
Table 5: Regression MSE $\times 10^{-3}$ (mean $\pm$ std) on the Pendulum dataset. We use a custom RoMAE model size.
Model Regression MSE ($\times 10^{-3}$)
@@
DESC ELAsTiCC Challenge
The DESC ELAsTiCC Challenge is a multi-variate irregular time-series dataset consisting of $\sim$1.8M simulated light curves and 36 classes of astronomical objects. Each light curve consists of 6 irregularly sampled channels (called ‘bands’). To embed this data in RoMAE, we follow the procedure described in Section 4.2. This results in a 2 dimensional positional embedding, where one dimension embeds the time, and the second embeds the channel index. Although ELAsTiCC has additional metadata for each light curve, we compare performance only across raw light curves. We train RoMAE-tiny by conducting full pre-training for 200 epochs with a masking ratio of 75%, then fine-tuning for 25 epochs. For full details and a visualization of the data we refer to Appendix D.6.
Table 5 shows our results using two sizes of RoMAE, and compares with ATAT [7], a Transformer architecture specialized for ELAsTiCC. Despite the latter’s specialization, RoMAE-tiny-shallow (with a comparable number of parameters as ATAT) improves over ATAT by about .08 F-score. The larger RoMAE-tiny achieves an improvement of .18 F-score. A key reason for RoMAE’s better performance might be that ATAT does not conduct any pre-training. In the case of RoMAE-tiny, the larger scale of the model also likely plays a role.
UEA Multivariate Time-series Archive
Model
Dataset TST [65] mTAN [46] S5 [51] ContiFormer [11] RoMAE-tiny
BM 0.9667 0.9917 0.9833 0.9750 0.9917
CT 0.9742 0.9529 0.9610 0.9833 0.9882
EP 0.9589 0.9203 0.9074 0.9324 0.9517
HB 0.7398 0.7789 0.7333 0.7561 0.7447
LSST 0.5520 0.5307 0.6389 0.6004 0.6225
We evaluate RoMAE on a variety of datasets from the UEA Multivariate Time-series Archive [3]. To make the datasets irregular we follow the procedure outlined by Kidger et. al. [30], dropping 30% of the observations. Because all variates are present at each time-step, the only irregular dimension is time. Therefore, to embed this data in RoMAE, we combine all variates per time-step into one embedding. This is a much simpler setup than the one used for the ELAsTiCC dataset in Section 5.4. For each dataset we conduct pre-training for 400 epochs. When fine-tuning, we found it necessary to change hyper-parameters between different datasets. For more details on our experimental setup we refer to Appendix D.3.
Table 6 presents the mean accuracy from 3 full training runs (pretraining + finetuning), as well as published results from a variety of models. Overall, RoMAE-tiny performs similarly or better than the comparators, and is able to handle the various datasets without issues. All the datasets trained on are relatively small, with some being on the order of hundreds of samples. Therefore, this experiment also shows how data-efficient RoMAE can be.
Irregular Time-series Regression: Pendulum Dataset
The Pendulum dataset [51] is an irregular time-series dataset consisting of irregularly sampled images of a pendulum. To embed the images in RoMAE, we use a patch size of $(1,24,24)$ for (time, height, width). This corresponds to 1 embedding per time-step/image. RoMAE is trained directly on regression without any pre-training, predicting the sine and cosine of the angle of the pendulum which follows a non-linear dynamical system. We use a custom size for RoMAE which contains only 2 layers and an MLP hidden size of 30. This is much smaller than RoMAE-tiny and provides for a fairer comparison with other models trained on this dataset. For additional information on the dataset and full experimental details we refer to Appendix D.4. After training on 20 different seeds, the mean MSE and standard deviation of RoMAE on the Pendulum dataset are reported in Table 5. RoMAE outperforms specialized Transformer based models such as ContiFormer [12], as well as state space models such as S5 [51].